OpenStudy (anonymous):

Solve for b. (sqrt[b^4/25] = 4)

OpenStudy (anonymous):

\[\sqrt{b ^{4}/25} = 4\]

OpenStudy (anonymous):

square on both sides..

OpenStudy (anonymous):

\[B ^{4}/25 = 4^2\]

OpenStudy (anonymous):

\[b ^{4}/25 = 16\]

OpenStudy (anonymous):

ok now simplify!

OpenStudy (anonymous):

\[\LARGE \sqrt{\frac{b^4}{25}}=4\] \[\LARGE \frac{\sqrt{b^4}}{\sqrt{25}}=4\] \[\LARGE \frac{b^2}{5}=4\] \[\LARGE b^2=20\] \[\LARGE b=?\]

OpenStudy (anonymous):

or you could do it that way :D..

OpenStudy (anonymous):

is the same :) ..

OpenStudy (anonymous):

find the sqrt of both sides. B and 4.47 B = 4.47?

OpenStudy (anonymous):

\[\LARGE b^2=20\] \[\LARGE b=\pm\sqrt{20}\] \[\LARGE b=\pm\sqrt{4\cdot 5}\] \[\LARGE b=\pm2\sqrt{5}\]

OpenStudy (anonymous):

A. 2 B. \[\pm5\sqrt{2}\] C. 3 D. \[\pm2\sqrt{5}\]